OpenStudy (anonymous):

sin2(thelta)/1+ cos2(thelta) =4

OpenStudy (anonymous):

\[\frac{ \sin ^{2}(x) }{ 1 }+\cos ^{2}(x) =4\] like this? but with x instead of theta?